# Cognitio Emergens: Agency, Dimensions, and Dynamics in Human–AI Knowledge Co-Creation

### v1

@@ Abstract
Scientific knowledge creation is fundamentally transforming as humans and AI systems evolve beyond tool-user relationships into co-evolutionary epistemic partnerships. When AlphaFold revolutionized protein structure prediction, researchers described engaging with an epistemic partner that reshaped how they conceptualized fundamental relationships. This article introduces Cognitio Emergens (CE), a framework addressing critical limitations in existing models that focus on static roles or narrow metrics while failing to capture how scientific understanding emerges through recursive human-AI interaction over time.
CE integrates three components addressing these limitations: Agency Configurations describing how authority distributes between humans and AI (Directed, Contributory, Partnership), with partnerships dynamically oscillating between configurations rather than following linear progression; Epistemic Dimensions capturing six specific capabilities emerging through collaboration across Discovery, Integration, and Projection axes, creating distinctive "capability signatures" that guide development; and Partnership Dynamics identifying forces shaping how these relationships evolve, particularly the risk of epistemic alienation where researchers lose interpretive control over knowledge they formally endorse.
Drawing from autopoiesis theory, social systems theory, and organizational modularity, CE reveals how knowledge co-creation emerges through continuous negotiation of roles, values, and organizational structures. The framework provides practical implementation guidance for researchers detecting dimensional imbalances, institutional leaders designing appropriate governance structures, and policymakers developing nuanced evaluation approaches. By reconceptualizing human-AI scientific collaboration as fundamentally co-evolutionary, CE offers a balanced perspective that neither uncritically celebrates nor unnecessarily fears AI’s evolving role, instead providing conceptual tools for cultivating partnerships that maintain meaningful human participation while enabling transformative scientific breakthroughs.
@@ 1. Introduction
Scientific knowledge creation is undergoing a fundamental transformation, shifting from human-centered processes augmented by tools to co-evolutionary partnerships between humans and artificial intelligence. When AlphaFold predicted protein structures with unprecedented accuracy in 2020, it marked more than a technical achievement, catalyzing a conceptual transformation in structural biology. Researchers described not merely using a tool but engaging with an epistemic partner that reshaped how they conceptualized structure-function relationships. Similar transformations now manifest across scientific domains: Google’s AI Co-Scientist autonomously generates testable hypotheses that researchers might never have formulated (Gottweis et al., 2025); SakanaAI systems produce peer-reviewed machine learning papers that introduce novel concepts rather than merely implementing human ideas (Yamada et al., 2025).
@@
These examples reveal a fundamental shift from AI as productivity enhancer to genuine epistemic partner in knowledge creation. This phenomenon—which we term Cognitio Emergens (from Latin: knowledge + emerging)—represents the dynamic emergence of scientific understanding through human-AI interaction that transcends what either could produce independently. This co-evolutionary process involves the mutual shaping of both human and artificial cognitive frameworks through sustained interaction. It yields knowledge structures neither would develop in isolation. Yet our conceptual frameworks remain anchored in outdated paradigms that position AI as either sophisticated tool (Amershi et al., 2019) or potential risk (Fragiadakis, Diou, Kousiouris, & Nikolaidou, 2025), failing to capture this emergent nature of human-AI epistemic partnerships.
@@
This conceptual gap creates profound tensions in scientific practice, hindering our ability to effectively cultivate and manage these powerful new relationships. Research communities now face unprecedented challenges: organizational structures struggle to adapt when knowledge creation transcends traditional human-centered epistemic communities (Liu & Shen, 2025; Lou, Lu, Raghu, & Zhang, 2025); researchers struggle to maintain interpretive ownership of insights emerging from interactions they cannot fully trace (Shanahan, 2025). Most fundamentally, we lack adequate theories for collaboration when boundaries between human and AI cognition increasingly blur.
@@
Existing frameworks inadequately address these tensions. Taxonomic approaches offer static classifications that miss the dynamic evolution of human-AI relationships (Dellermann et al., 2021; Holter & El-Assady, 2024). Metric-centered evaluations reduce complex epistemic processes to simplified performance indicators (Fragiadakis et al., 2025). Valuable work on epistemic shifts (Chen, 2025; Seeber et al., 2018) and organizational perspectives (Liu & Shen, 2025; Lou et al., 2025) remains fragmented, failing to integrate micro-interactions with macro-level knowledge structures. Even promising directions like Silver and Sutton’s (2025) “era of experience” (Silver & Sutton, 2025) remain predominantly agent-centric, neglecting the co-evolutionary nature of human-AI epistemic partnerships. Consequently, current approaches provide insufficient guidance for navigating the complexities of emergent knowledge co-creation.
@@
To address these critical limitations, we present the Cognitio Emergens (CE) framework—an integrated theoretical structure that reconceptualizes human-AI scientific collaboration as a co-evolutionary epistemic partnership with emergent properties distinct from either human or AI capabilities alone. The framework functions as an integrated system comprising three interrelated components that together capture the emergent nature of human-AI scientific collaboration:
@@
    Agency Configurations provide a dynamic model of how epistemic authority distributes between humans and AI, capturing the fluid transitions between Directed, Contributory, and Partnership modes that characterize mature collaborations.
@@
    Epistemic Dimensions identify six specific capabilities that emerge through sustained interaction, organized along Discovery, Integration, and Projection axes. These dimensions represent emergent properties of the partnership rather than attributes of either participant alone.
@@
    Partnership Dynamics illuminate the generative opportunities and potential vulnerabilities that shape how these relationships evolve, particularly the risk of epistemic alienation where researchers lose interpretive ownership of knowledge claims.
Together, these components form a comprehensive model for understanding and guiding how scientific knowledge emerges through human-AI interaction.
The CE framework’s distinctive theoretical contribution emerges from its integration of multiple traditions typically treated in isolation. Autopoiesis theory (Maturana & Varela, 1980) provides a foundation for understanding how recursive interaction between distinct cognitive systems generates emergent properties through structural coupling. Social systems theory (Luhmann, 1995) offers insights into how distinct communicative frameworks maintain boundaries while evolving through interaction. Organizational modularity theory (Baldwin & Clark, 2000) contributes understanding of how knowledge flows across boundaries and how structural design influences epistemic development. By synthesizing these perspectives, CE reconceptualizes human-AI scientific collaboration as a co-evolutionary process where capabilities emerge through recursive interaction. The framework directly addresses temporal myopia, static-agent assumptions, reductive success criteria, absence of value negotiation, and under-theorized organizational contexts that limit existing approaches, offering a more holistic understanding of these complex partnerships.
The significance of this work extends beyond theoretical contribution to address urgent practical needs in scientific practice. As AI systems increasingly transition from tools to partners in scientific knowledge creation, frameworks like Cognitio Emergens become essential for navigating this transformation. The approaches outlined here offer researchers, institutional leaders, and policymakers conceptual tools for cultivating partnerships that maximize transformative potential while preserving the epistemic integrity that underlies scientific progress. By understanding the evolutionary dynamics of these partnerships, we can foster scientific practices that harness AI’s computational capabilities while maintaining meaningful human participation in knowledge creation.
This paper proceeds by examining existing literature (Section 2), developing the CE framework (Section 3), providing implementation guidance (Section 4), and discussing implications and limitations (Section 5). Throughout, we maintain a balanced perspective that neither uncritically celebrates nor unnecessarily fears AI’s evolving role in scientific knowledge creation, instead offering a nuanced approach to understanding and guiding these emerging epistemic partnerships. As science advances into this new era, the future of knowledge creation depends not on human-only or AI-dominated processes, but on thoughtfully constructed partnerships that leverage the distinctive capabilities of both while acknowledging their fundamentally co-evolutionary nature.
@@ 2. Literature Review
This review examines frameworks for human–AI collaboration, identifying critical limitations that the Cognitio Emergens framework addresses. We organize the review around key approaches, highlighting persistent gaps concerning temporal dynamics, agent evolution, multidimensional assessment, epistemic values, and organizational context. This systematic analysis provides the foundation for the integrated framework introduced in the previous section.
@@ 2.1 From Static Heuristics to Dynamic Taxonomies
Early human-AI frameworks relied on static interaction principles that failed to account for the dynamic, evolving nature of advanced AI systems. These early approaches embodied a static-agent assumption, rendering them incapable of accounting for evolving AI capabilities and the dynamic nature of human-AI interaction. This limitation spurred the development of new frameworks (Amershi et al., 2019).
More recent taxonomies attempt to address this dynamism. For example, Dellermann et al. (2021) classify hybrid intelligence systems using dimensions like task characteristics and interaction modes. Holter & El-Assady (2024) deconstruct mixed-initiative systems along agency, interaction, and adaptation dimensions. Others classify AI roles based on initiation spectrum, intelligence scope, and cognitive mode (Song, Zhu, & Luo, 2024). While acknowledging adaptation, these taxonomies often model roles as discrete configurations, potentially overlooking the continuous co-evolution central to mature partnerships (Holter & El-Assady, 2024). Furthermore, most taxonomies inadequately address multi-loop learning processes—recursive feedback cycles (Argyris, 1976; Tosey, Visser, & Saunders, 2012) where both epistemic frameworks and learning procedures co-evolve through sustained interaction (Lou et al., 2025). Thus, even dynamic taxonomies often fall short of capturing the deep, reciprocal evolution characteristic of advanced human-AI collaboration.
While taxonomic approaches attempted to capture more dynamic aspects of collaboration, their categorical nature prompted researchers to explore more quantitative evaluation methods.
@@ 2.2 Metric-Centric Evaluation Frameworks
Attempting to overcome taxonomic limitations, metric-centric frameworks quantify collaboration through performance indicators, yet introduce their own significant constraints. Empirical studies applying such metrics show promising results; for instance, Woelfle et al. (2024) demonstrate that human–LLM teams can outperform individuals in specific evidence appraisal tasks.
However, these metric-based approaches introduce their own significant limitations. They often exhibit temporal myopia, capturing only single-session snapshots rather than longitudinal capability evolution (Fragiadakis et al., 2025). They also suffer from scalar reductivism, collapsing complex, multidimensional epistemic labor into simplified scores. Consequently, the crucial relational dimension of human–AI collaboration—how the partnership itself develops and influences outcomes—remains undertheorized and largely unmeasured (Watkins, Moss, Raffa, & Nachman, 2025), limiting the insights gained from purely metric-driven evaluations.
@@ 2.3 Epistemic-Shift and Organizational Perspectives
Epistemic-shift and organizational perspectives expand beyond individual agents to examine how AI integration fundamentally transforms knowledge structures and institutional practices. Studies document how teams reorganize when AI becomes a teammate (Seeber et al., 2018) and how generative-AI infrastructures redistribute authority in specific contexts like education (Chen, 2025). These studies highlight the under-theorized value negotiation in many frameworks, revealing a neglect of how epistemic values themselves evolve through human–AI interaction.
Recent analyses (Lou et al., 2025) further expose a theoretical blind spot regarding how power dynamics and governance structures shape human–AI collaboration. This gap extends to regulatory and institutional constraints (compliance, liability) that fundamentally shape viable collaborative configurations. Many frameworks also fail to fully embrace a sociotechnical systems perspective, which would better articulate the mutual constitution of technological capabilities and social structures through iterative interaction (Lou et al., 2025; Watkins et al., 2025). Neglecting these organizational and socio-technical factors prevents a full understanding of how partnerships function in real-world contexts.
From an organizational design perspective, Baldwin and Clark’s (Baldwin & Clark, 2000) work on modular architectures offers relevant insights, showing how design choices distribute knowledge across components. Similarly, research on structural design (Joseph & Sengul, 2024) and ambidextrous organizations (O’Reilly & Tushman, 2013) highlights how architectural choices influence knowledge flows and the balance between exploration and exploitation. While valuable, these organizational perspectives often remain disconnected from the micro-level dynamics of human-AI interaction and epistemic emergence.
These organizational insights, while valuable, still largely operate within episodic rather than continuous evolutionary frameworks.
@@ 2.4 Toward Continuous Experience: An Emerging Direction
Seeking to address the limitations of static and short-term views, more recent approaches emphasize continuous learning and adaptation, though significant gaps remain. Silver and Sutton’s (Silver & Sutton, 2025) era-of-experience proposal represents an important advancement. They suggest next-generation AI agents will learn through continuous experience, using self-generated world models to maximize reward. This perspective rightly acknowledges the dynamic, ongoing nature of AI development often missed by static taxonomies.
However, this perspective remains predominantly agent-centric. It offers limited insight into how humans negotiate reward meaning in collaborative contexts or how organizational structures scaffold these continuous learning processes. Most significantly, it does not adequately theorize the co-evolutionary nature of human–AI epistemic partnerships—the crucial process by which human and machine cognition reciprocally reshape each other through structural coupling (Silver & Sutton, 2025; Watkins et al., 2025). Focusing primarily on the AI agent’s experience risks overlooking the emergent properties of the partnership itself.
@@ 2.5 Theoretical Lineage of Cognitio Emergens
The Cognitio Emergens framework bridges several theoretical traditions. It draws on autopoiesis (Maturana & Varela, 1980) to explain how feedback loops yield emergent behaviors in coupled systems. Luhmann’s social systems theory (Luhmann, 1995) provides insights into how AI and human researchers act as structurally coupled yet operationally distinct cognitive frameworks.
Shanahan’s work on emergent cognition (Shanahan, 2025) cautions that adaptive AI systems may approach “the void of inscrutability” where traditional interpretability frameworks collapse. Baldwin and Clark’s organizational modularity theory (Baldwin & Clark, 2000) reveals how well-defined interfaces between cognitive modules can nurture innovation while warning against rigid compartmentalization.
@@ 2.6 Critical Gaps and the Cognitio Emergens Framework
This synthesis reveals five persistent limitations across existing approaches: temporal myopia, evident in short-term evaluation studies (Fragiadakis et al., 2025; Liu & Shen, 2025); static-agent assumptions portraying roles as fixed rather than co-evolving (Holter & El-Assady, 2024; Watkins et al., 2025); reductive success criteria collapsing multidimensional knowledge work into scalar metrics (Fragiadakis et al., 2025; Lou et al., 2025); a limited understanding of AI’s impact on epistemic values, often neglecting how collaboration reshapes or potentially undermines human values and agency (Chen, 2025); and an under-theorized organizational context overlooking power, governance, and regulatory forces (Liu & Shen, 2025; Lou et al., 2025).
Table 1 systematically maps how existing approaches address the five critical dimensions identified in our analysis, highlighting where the CE framework offers distinctive contributions.
Table 1: Comparison of Frameworks Across Key Dimensions
Approach (Key Sources) Temporal Dynamics Agent Evolution Multi-dimensional Assessment Treatment of Epistemic Values Organizational Context
Interaction Heuristics (Amershi et al., 2019) Static snapshots Fixed capabilities Performance metrics Implicit values Minimal consideration
Dynamic Taxonomies (Dellermann et al., 2021; Holter & El-Assady, 2024) Limited adaptation Categorical evolution Multi-factor assessment Implicit values Limited integration
Metric Frameworks (Fragiadakis et al., 2025) Short-term measures Performance adaptation Scalar indicators Fixed criteria Background factor
Epistemic-Shift Studies (Chen, 2025; Seeber et al., 2018; Watkins et al., 2025) Episodic analysis Limited co-evolution Qualitative dimensions Explicit examination Contextual analysis
Continuous Experience (Silver & Sutton, 2025) Continuous learning Agent-centric evolution Reward-based metrics Limited negotiation Minimal consideration
Cognitio Emergens Evolutionary trajectories Co-evolutionary development Six-dimension framework Value co-construction Integrated perspective
The Cognitio Emergens framework addresses these five gaps simultaneously, offering an integrated theoretical scaffold that reconceptualizes human–AI scientific collaboration as fundamentally co-evolutionary. By capturing both the structural configurations and emergent properties of these partnerships, CE provides a more comprehensive basis for understanding and guiding the future of scientific knowledge creation.

### v2

# Ginwidth=keepaspectratio

@@ Abstract
Human-AI scientific collaboration has evolved from tool-user relationships into co-evolutionary partnerships. When AlphaFold improved protein structure prediction, researchers engaged with an epistemic partner that transformed their approach to structure-function problems. Yet existing frameworks position AI as either sophisticated tool or potential risk, overlooking how scientific understanding emerges through recursive interaction. We introduce Cognitio Emergens (CE), a framework that captures the co-evolutionary nature of human-AI epistemic partnerships.
Drawing from autopoiesis theory, social systems theory, and organizational modularity, CE integrates three components: Agency Configurations modeling how authority distributes through Directed, Contributory, and Partnership modes, with partnerships oscillating dynamically rather than following linear progression; Epistemic Dimensions capturing six capabilities along Discovery, Integration, and Projection axes, creating distinctive “capability signatures” that guide strategic development; and Partnership Dynamics identifying evolutionary forces including epistemic alienation—where researchers lose interpretive control over knowledge they formally endorse.
The framework equips researchers to diagnose dimensional imbalances, institutional leaders to design governance structures supporting multiple agency configurations, and policymakers to develop evaluations beyond simple performance metrics. By reconceptualizing human-AI collaboration as fundamentally co-evolutionary, CE provides conceptual tools for cultivating partnerships that preserve epistemic integrity while enabling transformative breakthroughs neither humans nor AI could achieve independently.
@@ 1. Introduction
Scientific knowledge creation is shifting from human-centered processes augmented by tools to partnerships between humans and artificial intelligence. When AlphaFold improved protein structure prediction in 2020, researchers described engaging with an epistemic partner that changed how they approached structure-function relationships. This pattern appears across domains: Google’s AI Co-Scientist autonomously generates testable hypotheses (Gottweis et al. 2025); SakanaAI systems produce peer-reviewed machine learning papers introducing novel concepts (Yamada et al. 2025).
@@
These examples reveal AI functioning as an epistemic partner in knowledge creation. We term this Cognitio Emergens—the emergence of scientific understanding through human-AI interaction that exceeds what either could produce independently. This process involves mutual shaping of human and artificial cognitive frameworks through sustained interaction, yielding knowledge structures neither would develop alone. Yet existing frameworks position AI as either sophisticated tool (Amershi et al. 2019) or potential risk (Fragiadakis et al. 2025), overlooking how these partnerships actually function.
@@
Research communities now face a fundamental challenge: existing theories cannot account for collaboration where boundaries between human and AI cognition blur. This gap between practice and theory creates tensions throughout scientific practice. Organizational structures designed for human researchers struggle to accommodate human-AI teams (Lou et al. 2025; Liu and Shen 2025). Researchers cannot always explain insights emerging from interactions they cannot fully trace (Shanahan 2025).
@@
Existing frameworks address these tensions inadequately. Taxonomic approaches offer static classifications that miss how human-AI relationships evolve (Dellermann et al. 2021; Holter and El-Assady 2024). Metric-centered evaluations reduce epistemic processes to performance indicators (Fragiadakis et al. 2025). Work on epistemic shifts (Seeber et al. 2018; Chen 2025) and organizational perspectives (Lou et al. 2025; Liu and Shen 2025) remains fragmented, failing to integrate micro-interactions with macro-level structures. Even Silver and Sutton’s (2025) “era of experience” (Silver and Sutton 2025) remains agent-centric, neglecting co-evolutionary aspects of human-AI partnerships. Current approaches provide insufficient guidance for emergent knowledge co-creation.
@@
To address these limitations, we present the Cognitio Emergens (CE) framework—an integrated structure for understanding human-AI scientific collaboration as co-evolutionary partnership with emergent properties. The framework comprises three interrelated components:
@@
    Agency Configurations model how epistemic authority distributes between humans and AI, capturing transitions between Directed, Contributory, and Partnership modes.
@@
    Epistemic Dimensions identify six capabilities that emerge through sustained interaction, organized along Discovery, Integration, and Projection axes.
@@
    Partnership Dynamics reveal opportunities and vulnerabilities shaping relationship evolution, including epistemic alienation where researchers lose interpretive ownership of knowledge claims.
Together, these components provide a model for understanding how scientific knowledge emerges through human-AI interaction.
The CE framework integrates multiple traditions. Autopoiesis theory (Maturana and Varela 1980) explains how recursive interaction between cognitive systems generates emergent properties through structural coupling. Social systems theory (Luhmann 1995) shows how distinct communicative frameworks maintain boundaries while evolving through interaction. Organizational modularity theory (Baldwin and Clark 2000) reveals how structural design influences knowledge flows and epistemic development. Synthesizing these perspectives, CE treats human-AI collaboration as co-evolutionary process where capabilities emerge through recursive interaction, addressing temporal myopia, static-agent assumptions, reductive success criteria, absence of value negotiation, and under-theorized organizational contexts in existing approaches.
Beyond theoretical synthesis, the framework addresses practical needs in scientific practice. As AI systems transition from tools to partners in knowledge creation, frameworks like Cognitio Emergens help navigate this transformation. The framework provides researchers, institutional leaders, and policymakers with tools for cultivating partnerships that preserve epistemic integrity while enabling scientific progress. Understanding partnership dynamics enables practices that integrate the distinctive capabilities of both humans and AI.
This paper examines existing literature (Section 2), develops the CE framework (Section 3), provides implementation guidance (Section 4), and discusses implications and limitations (Section 5). The future of knowledge creation depends on partnerships that leverage the distinctive capabilities of both humans and AI while acknowledging their co-evolutionary nature.
@@ 2. Literature Review
This review examines frameworks for human–AI collaboration, identifying limitations the Cognitio Emergens framework addresses. The review highlights gaps concerning temporal dynamics, agent evolution, multidimensional assessment, epistemic values, and organizational context.
@@ 2.1 From Static Heuristics to Dynamic Taxonomies
Early human-AI frameworks assumed static agent capabilities (Amershi et al. 2019), overlooking how AI systems evolve during use and how interaction patterns shift as agents learn. Recent taxonomies address this limitation: Dellermann et al. (2021) classify hybrid intelligence systems using task characteristics and interaction modes; Holter and El-Assady (2024) analyze mixed-initiative systems along agency, interaction, and adaptation dimensions; Song, Zhu, and Luo (2024) classify AI roles by initiation spectrum, intelligence scope, and cognitive mode. However, these taxonomies model collaboration as discrete configurations rather than continuous co-evolution. Most overlook multi-loop learning processes—recursive cycles where both agents’ epistemic frameworks and learning procedures change through sustained interaction (Argyris 1976; Tosey, Visser, and Saunders 2012; Lou et al. 2025).
@@ 2.2 Metric-Centric Evaluation Frameworks
Metric-centric frameworks quantify collaboration through performance indicators. Empirical studies show promise: Woelfle et al. (2024) demonstrate that human–LLM teams outperform individuals in specific evidence appraisal tasks.
However, metric-based approaches have limitations. They exhibit temporal myopia, capturing single-session snapshots rather than longitudinal evolution (Fragiadakis et al. 2025). They suffer from scalar reductivism, collapsing complex epistemic labor into simplified scores. The relational dimension—how the partnership develops and influences outcomes—remains undertheorized and unmeasured (Watkins et al. 2025).
@@ 2.3 Epistemic-Shift and Organizational Perspectives
Epistemic-shift and organizational perspectives examine how AI integration transforms knowledge structures and institutional practices. Studies document how teams reorganize when AI becomes a teammate (Seeber et al. 2018) and how generative-AI infrastructures redistribute authority (Chen 2025). These studies reveal neglect of how epistemic values evolve through human–AI interaction.
Recent analyses (Lou et al. 2025) expose how power dynamics and governance structures shape human–AI collaboration. Regulatory and institutional constraints (compliance, liability) shape viable collaborative configurations. Many frameworks fail to embrace a sociotechnical systems perspective that articulates mutual constitution of technological capabilities and social structures (Lou et al. 2025; Watkins et al. 2025).
From an organizational design perspective, Baldwin and Clark (2000) show how design choices distribute knowledge across components. Research on structural design (Joseph and Sengul 2024) and ambidextrous organizations (O’Reilly and Tushman 2013) reveals how architectural choices influence knowledge flows and balance between exploration and exploitation. These organizational perspectives often remain disconnected from micro-level dynamics of human-AI interaction.
@@ 2.4 Toward Continuous Experience: An Emerging Direction
Recent approaches emphasize continuous learning and adaptation. Silver and Sutton’s (Silver and Sutton 2025) era-of-experience proposal suggests next-generation AI agents will learn through continuous experience, using self-generated world models to maximize reward. This acknowledges the dynamic nature of AI development missed by static taxonomies.
However, this perspective remains agent-centric. It offers limited insight into how humans negotiate reward meaning collaboratively or how organizational structures scaffold continuous learning. It does not adequately theorize the co-evolutionary nature of human–AI partnerships—how human and machine cognition reciprocally reshape each other through structural coupling (Silver and Sutton 2025; Watkins et al. 2025). Focusing on the AI agent’s experience overlooks emergent properties of the partnership.
@@ 2.5 Theoretical Lineage of Cognitio Emergens
The Cognitio Emergens framework bridges several traditions. Autopoiesis (Maturana and Varela 1980) explains how feedback loops yield emergent behaviors in coupled systems. Social systems theory (Luhmann 1995) shows how AI and human researchers act as structurally coupled yet operationally distinct cognitive frameworks. Shanahan’s work on emergent cognition (Shanahan 2025) cautions that adaptive AI may approach “the void of inscrutability” where interpretability frameworks collapse. Organizational modularity theory (Baldwin and Clark 2000) reveals how well-defined interfaces between cognitive modules nurture innovation.
@@ 2.6 Critical Gaps and the Cognitio Emergens Framework
This synthesis reveals five limitations: temporal myopia in short-term evaluations (Fragiadakis et al. 2025; Liu and Shen 2025); static-agent assumptions portraying roles as fixed rather than co-evolving (Holter and El-Assady 2024; Watkins et al. 2025); reductive success criteria collapsing multidimensional knowledge work into scalar metrics (Fragiadakis et al. 2025; Lou et al. 2025); limited understanding of AI’s impact on epistemic values (Chen 2025); and under-theorized organizational context overlooking power, governance, and regulatory forces (Lou et al. 2025; Liu and Shen 2025).
Table 1 systematically maps how existing approaches address the five critical dimensions identified in our analysis, highlighting where the CE framework provides distinctive contributions.
Approach (Key Sources) Temporal Dynamics Agent Evolution Multidimensional Assessment Treatment of Epistemic Values Organizational Context
Interaction Heuristics (Amershi et al. 2019) Static snapshots Fixed capabilities Performance metrics Implicit values Minimal consideration
Dynamic Taxonomies (Dellermann et al. 2021; Holter and El-Assady 2024) Limited adaptation Categorical evolution Multi‑factor assessment Implicit values Limited integration
Metric Frameworks (Fragiadakis et al. 2025) Short‑term measures Performance adaptation Scalar indicators Fixed criteria Background factor
Epistemic‑Shift Studies (Seeber et al. 2018; Chen 2025; Watkins et al. 2025) Episodic analysis Limited co‑evolution Qualitative dimensions Explicit examination Contextual analysis
Continuous Experience (Silver and Sutton 2025) Continuous learning Agent‑centric evolution Reward‑based metrics Limited negotiation Minimal consideration
Cognitio Emergens Evolutionary trajectories Co‑evolutionary development Six‑dimension framework Value co‑construction Integrated perspective
The Cognitio Emergens framework addresses these five gaps, offering a scaffold that treats human–AI collaboration as co-evolutionary. By capturing structural configurations and emergent properties, CE provides a basis for understanding scientific knowledge creation in human-AI partnerships.